# No.2 Linggong Road, Dalian, Liaoning, 116024, P.R.China

### v3

@@ 3.2 Baseline models and training
All models involved (BDT, MLP, LCF in two strategies) are trained using the same configuration: binary cross entropy as the loss function, Adam as the optimizer with a learning rate 0.001, a batch size of 512, and 200 training epochs. Early stopping is not applied as training of LCF exhibits plateau phases in loss. The centers for the LCF model are set as -2, 2, 0, 0, -1, -3, -5, 5, -1.8, -1.4 for each mock feature as shown in figure 2 and 80, 0.15, 0.025, 2, 2, 0.3 for each observable in the diboson dataset as shown in figure 4. Each center is chosen to divide the respective observable’s distribution into two regions, enabling the LCF to learn boundaries on one side (for left or right cases) or both sides (for middle or edge cases). The choice of center values is not strict. For single-sided cuts, the center can be placed anywhere except at the expected boundary. For double-sided cuts, it is ideally positioned near the middle of the distribution to enable learning of boundaries on both sides. If misplaced—e.g., at the distribution’s leftmost edge—the double-sided cuts will degenerate into single-sided cuts, mismatching the distribution and reducing performance. Models are trained continuously ten times to demonstrate robustness and stability.
@@ 3.3 Performance evaluation
  • Significance: A measure of sensitivity to a rare signal against a dominant background in high-energy physics. It’s computed as $S/\sqrt{B}$, where $S=\epsilon_{s}\times\sigma_{s}\times\mathcal{L}$ is the expected number of signal events and $B=\epsilon_{b}\times\sigma_{b}\times\mathcal{L}$ is the expected number of background events after all learned cuts. Here, $\epsilon_{s}=$TP / total test signal events is the signal efficiency, $\epsilon_{b}=$FP / total test background events is the background efficiency. The integrated luminosity $\mathcal{L}$ is assumed to be 3000 $\text{fb}^{-1}$. For the mock datasets, we assume: $\sigma_{s}=1$ pb (rare signal) and $\sigma_{b}=10^{6}$ pb (abundant background). For the real diboson dataset, we estimate cross sections based on the original processes: approximately 0.7644 pb for the diboson signal and $1.806\times 10^{5}$ pb for the QCD background. Since ref. 1603.09349 does not provide cross sections after their full processing pipeline, we derived cross sections independently using only anti-kt jet clustering ($R=1.2$). These nominal cross sections are used for basic significance comparisons and differ from those in the reference due to the omitted trimming and selection steps.
@@ 4.3 Robustness to redundancy
Table 4 shows that both strategies achieve almost the same performance. Across ten training runs, $x_{7}$’s cuts in the parallel strategy show inconsistent importance scores relative to the baseline—while eight runs yield scores below the baseline, two runs obtain scores above it (0.0183 and 0.0187). When $x_{7}$’s importance exceeds the baseline, the cut on it is applied during inference; when below, it is ignored entirely. Such treatment leads to higher metric errors in the parallel strategy then those in the sequential strategy: accuracy error of $6.8\%$ vs $0.0\%$ and significance error of $0.5$ vs $0.0$. Despite this variability, the average importance of $x_{7}$ is similar between both strategies, leading to the final similarity in significance obtained with the two strategies.
These findings highlight the LCF models’ robustness in detecting and effectively ignoring features with negligible separation power.
@@ 4.6 Benchmark on diboson classification
The training time of the LCF model is at the same level as the BDT and MLP models, which is lightweight to train compared to modern deep neural networks. The significance of the LCF models, however, is slightly worse than the BDT and MLP models because they are simpler and shallower networks essentially. The parallel strategy this time doesn’t outperform the sequential strategy in terms of significance. The parallel strategy’s more conservative cut boundaries, while effective in the Mock datasets, here incur signal efficiency costs that outweigh the background reduction benefits in this diboson dataset. This demonstrates the fundamental trade-off inherent in the parallel strategy’s approach. From this perspective, the sequential strategy is always more stable (lower variance of metrics across training runs) and more accurate (as a classifier with higher accuracy). However, the parallel strategy gains advantages due to the cuts that achieve lower classification accuracy—it always filters out more events. Combined with results from Mock datasets, such inaccuracy doesn’t lead to a consistent significance improvement. The sequential strategy is a better choice as a classifier, just like all other neural networks that are trained to reduce the loss designed for classification. On the other side, the parallel strategy tends to achieve better significance values in most cases, though the optimal choice depends on the specific dataset characteristics. Both strategies can be quickly validated in LCF and the better one can be chosen.

### v4

@@ 3.2 Baseline models and training
All models involved (BDT, MLP, LCF in two strategies) are trained using the same configuration: binary cross entropy as the loss function, Adam as the optimizer with a learning rate 0.001, a batch size of 512, and 200 training epochs. Early stopping is not applied as training of LCF exhibits plateau phases in loss. The centers for the LCF model are set as -2, 2, 0, 0, -1, -3, -5, 5, -1.8, -1.4 for each mock feature, as shown in figure 2, and 80, 0.15, 0.025, 2, 2, 0.3 for each observable in the diboson dataset, as shown in figure 4. Each center is chosen to divide the respective observable’s distribution into two regions, enabling the LCF to learn boundaries on one side (for left or right cases) or both sides (for middle or edge cases). Models are trained continuously ten times to demonstrate robustness and stability.
@@ 3.3 Performance evaluation
  • Significance: A measure of sensitivity to a rare signal against a dominant background in high-energy physics. It’s computed as $S/\sqrt{B}$, where $S=\epsilon_{s}\times\sigma_{s}\times\mathcal{L}$ is the expected number of signal events and $B=\epsilon_{b}\times\sigma_{b}\times\mathcal{L}$ is the expected number of background events after all learned cuts. Here, $\epsilon_{s}=$TP / total test signal events is the signal efficiency, $\epsilon_{b}=$FP / total test background events is the background efficiency. The integrated luminosity $\mathcal{L}$ is assumed to be 3000 $\text{fb}^{-1}$. For the mock datasets, we assume: $\sigma_{s}=1$ pb (rare signal) and $\sigma_{b}=10^{6}$ pb (abundant background). For the real diboson dataset, we estimate the cross sections based on the original processes: approximately 0.7644 pb for the diboson signal and $1.806\times 10^{5}$ pb for the QCD background. Since ref. 1603.09349 does not provide the cross sections after their full processing pipeline, we derived them independently using only anti-kt jet clustering ($R=1.2$). These nominal cross sections are used for basic significance comparisons and differ from those in the reference due to the omitted trimming and selection steps.
@@ 4.3 Robustness to redundancy
Table 4 shows that both strategies achieve almost the same performance. This finding highlights the LCF models’ robustness in detecting and effectively ignoring features with negligible separation power.
@@ 4.6 Benchmark on diboson classification
The training time of the LCF model is at the same level as the BDT and MLP models, which is lightweight to train compared to modern deep neural networks. The significance of the LCF models, however, is slightly worse than the BDT and MLP models because they are simpler and shallower networks essentially. The parallel strategy this time doesn’t outperform the sequential strategy in terms of significance. The parallel strategy’s more conservative cut boundaries, while effective in the Mock datasets, here incur signal efficiency costs that outweigh the background reduction benefits in this diboson dataset. This demonstrates the fundamental trade-off inherent in the parallel strategy’s approach. From this perspective, the sequential strategy is always more stable (lower variance of metrics across training runs) and more accurate. However, the parallel strategy gains advantages due to the cuts that achieve lower classification accuracy—it always filters out more events. Combined with results from Mock datasets, such inaccuracy doesn’t lead to a consistent significance improvement. The sequential strategy is a better choice as a classifier, just like all other neural networks that are trained to reduce the loss designed for classification. On the other hand, the parallel strategy tends to achieve better significance values in most cases, although the optimal choice depends on the specific dataset characteristics. Both strategies can be quickly validated in LCF and the better one can be chosen.